# Spin-free orbital entropy, mutual information, and entanglement analysis

### v1

@@ 5 Conclusions
We have defined spin-free analogues of orbital entropy, pair entropy, and mutual information and investigated their properties. We have shown that the spin-free counterparts cannot be bigger than the original spin-including quantities. Typically they are smaller, which indicates that a part of the strong correlation is actually a “static” correlation due to spin couplings. We have shown and numerically verified the invariance of the spin-free entropy and mutual information with respect to the $M_{s}$ component of the investigated state and have observed that the spin-free quantities are close to the original ones for the maximum $M_{s}$ component of the high-spin case. We studied two realistic examples of complexes with iron-sulfur bonds which exhibit strong correlation and by comparison of the spin-free and spin-including correlation measures we were able to identify to which extent the correlation is due to spin couplings. A trend of increase of the total spin-free entropy with decreasing spin of the state has been observed. Since the computation of the spin-free entanglement analysis does not require any significant additional computational cost, we think that it can be routinely used as a useful supplement to the analysis using the “traditional” spin-dependent orbital entropy and mutual information.

### v2

@@ 5 Conclusions
We have defined spin-free analogues of orbital entropy, pair entropy, and mutual information and investigated their properties. We have shown that the spin-free counterparts cannot be bigger than the original spin-including quantities. Typically they are smaller, which indicates that a part of the strong correlation is actually a “static” correlation due to spin couplings. We have shown and numerically verified the invariance of the spin-free entropy and mutual information with respect to the $M_{s}$ component of the investigated state and have observed that the spin-free quantities are close to the original ones for the maximum $M_{s}$ component of the high-spin case. We studied two realistic examples of complexes with iron-sulfur bonds which exhibit strong correlation and by comparison of the spin-free and spin-including correlation measures we were able to identify to which extent the correlation is due to spin couplings. A trend of increase of the total spin-free entropy with decreasing spin of the state has been observed. Since the computation of the spin-free correlation analysis does not require any significant additional computational cost, we think that it can be routinely used as a useful supplement to the analysis using the “traditional” spin-dependent orbital entropy and mutual information.